# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

### v3

@@ 5 When will Personalization Improve Communication Efficiency
(ii) # gradient $\begin{split}&\text{(ii) \# gradient evaluations}=\tilde{\mathcal{O}}\left(\kappa\cdot\log\frac{1}{\varepsilon}\right).\end{split}$ (14)
$\bullet$ Communication cost. Corollary 1 shows as $\lambda$ increases from 00 to $\infty$, the communication complexity increases from $\tilde{\mathcal{O}}(1)$ to $\tilde{\mathcal{O}}(\kappa)$. The result reveals a higher personalization degree requires less communication cost to collaboratively learn the models. In the full personalization limit with $\lambda\rightarrow 0$, the order of the communication cost reduces to near constant $\tilde{\mathcal{O}}(1)$, which corresponds to pure local training. It is worth noting that even personalization can reduce communication complexity, as our ultimate goal is to accurately estimate each model so we can not simply pursue higher communication efficiency.
$\bullet$ Computation cost. In contrast to the communication complexity, the gradient evaluations per round $K$ decrease with $\lambda$ increasing. Therefore, a higher degree of personalization requires a more intensive update of the local models once receiving a new global model ${\bm{w}}^{(g)}$, so as to adapt to each client’s local data set. Note that although a small $\lambda$ requires a larger $K$, the overall computation cost given by (LABEL:eq:comp-cost) is independent of $\lambda$. As such, model personalization (small $\lambda$) can provably reduce the communication cost without any extra computation overhead.
[b] Methods Convexity Bounded Domain Computation Cost per Round Communication Cost FedProx (Li et al., 2020a) cvx (1) $\mathcal{O}(\Delta/(\rho\varepsilon))$ pFedMe (T Dinh et al., 2020) s-cvx (2) $\mathcal{O}(1/{(mR\varepsilon)})$ Ditto (Li et al., 2021) s-cvx $\mathcal{O}(1)$ $\mathcal{O}(1/\varepsilon)$ L2GD (Hanzely & Richtárik, 2020) s-cvx $\mathcal{O}(1+\frac{L}{\lambda})$ $\mathcal{O}({\frac{2\lambda\kappa}{\lambda+L}\log(\frac{1}{\varepsilon})})$ Algorithm 3 (Chen et al., 2023b) s-cvx $\mathcal{O}(\lambda/\varepsilon)$ $\mathcal{O}((\lambda\vee 1)/\varepsilon)$ Our Result s-cvx $\mathcal{O}(\frac{\lambda+L}{\lambda+\mu})$ $\mathcal{O}(\kappa\frac{\lambda+\mu}{\lambda+L}\log{\frac{1}{\varepsilon}})$
Table 2: Comparison of the communication and computation complexity in existing works of analyzing the objective in (1) (cvx denotes convexity and s-cvx denotes strong convexity).
@@
We end this section with a comparison between our results and the existing algorithms for solving FedProx, as reported in Table 2. Our result shows how the personalization degree trades off the communication cost and local computation cost, whereas most of the existing works either don’t explicitly analyze the local computational cost (Li et al., 2020a; T Dinh et al., 2020), or fail to demonstrate such a trade-off (Chen et al., 2023b). Hanzely & Richtárik (2020) also reveals similar tradeoffs, but their work is a randomized method that requires more coordination between server and clients, and more importantly, when $\lambda$ is small, the computation cost per round is unbounded.
6 Simulation
In this section, we provide empirical validation for our theoretical results, evaluating FedProx on both synthetic and real datasets with convex and non-convex loss functions. We first examine the effect of personalization on statistical accuracy under different levels of heterogeneity (Section 6.2), followed by an analysis of its impact on communication efficiency (Section 6.3). Detailed descriptions of the experimental setup are available in Section 6.1 and Appendix C, and the complete anonymized codebase is accessible at https://anonymous.4open.science/r/fedprox-bilevel.
@@ 6.1 Experimental Details
Synthetic Dataset. To generate the synthetic data, we follow a similar procedure to prior works (Li et al., 2020a) but with some modifications to align with the setup in this paper. Specifically, for each client we generate samples $(\bm{X}_{k},\bm{y}_{k})$, where the labels $\bm{y}_{k}$ are produced by a logistic regression model $\bm{y}_{k}=\text{argmax}(\sigma({\bm{w}}_{k}^{\top}\bm{X}_{k}))$, with $\sigma$ being the sigmoid function. The feature vectors $\bm{X}_{k}$ are drawn from a multivariate normal distribution $\mathcal{N}(\bm{v}_{k},\bm{\Sigma})$, where $v_{k}\sim\mathcal{N}(0,1)$ and the covariance matrix $\bm{\Sigma}$ follows a diagonal structure $\bm{\Sigma}_{j,j}=j^{-1.2}$. The heterogeneity is introduced by sampling the model weights ${\bm{w}}_{k}$ for each client from a normal distribution $\mathcal{N}(0,R)$, where $R\in[0,3]$ controls the degree of variation (statistical heterogeneity) across clients’ data distributions.
Real Dataset. We use the MNIST, EMNIST, and CIFAR10 datasets for the real dataset analysis. We first formulate the problem as a convex classification task using multinomial logistic regression, then test the generalization of our theory to a more general, non-convex scenario by conducting experiments using different CNN classifiers. Following Li et al. (2020a), to impose statistical heterogeneity, we distribute the data across clients in the way that each client only has access to a fixed number of classes. The fewer classes each client has access to, the higher the statistical heterogeneity. The detail of data processing to set data heterogeneity for each dataset can be found in Appendix C.
@@
Implementation and Evaluation. For each setup, we evaluate the FedProx algorithm with three different choices of $\lambda$: small, medium, and large, with specific values varying based on the dataset (details provided in Appendix C). For the synthetic dataset, in alignment with theorem 2, we implement Algorithm 1 using a global step size $\gamma=(\lambda+L)/(\lambda L)$ and a local step size $\eta=(L+\lambda)^{-1}$. For the real dataset, since $L$ is unknown, we implement the same algorithm with a global step size $\gamma$ set to $1/(\lambda)$, while the local step size is determined via grid search. In terms of evaluation, for the synthetic dataset, we evaluate by tracking the ground-truth models and measuring error as the distance from these true models. For the real dataset, we report training loss and testing accuracy. Additional details, including hyperparameter setting and evaluation metric definition, can be found in Appendix C.
Figure 1: Statistical and total error of different methods. In the top row, subfigures (a), (b), and (c) represent statistical error, showing the difference between the minimizer and the underlying true model for Synthetic-Logit (logistic regression on synthetic data), MNIST-Logit (logistic regression on MNIST), and MNIST-CNN (CNN on MNIST), respectively. In the bottom row, subfigures (d), (e), and (f) represent total error, showing the error between the model at each round and the true model for the same datasets. (1) FedProx adapts between GlobalTrain and LocalTrain, achieving low error across varying levels of heterogeneity with a properly chosen $\lambda$; (2) Larger $\lambda$ (less personalization) requires more communication to converge, and vice versa.
6.2 Statistical Aspect: Balancing Local and Global Learning
We now explore how FedProx balances between GlobalTrain and LocalTrain by achieving different levels of personalization. For comparison, in addition to implementing FedProx with different choices of $\lambda$, we also implemented LocalTrain and GlobalTrain, with details specified in Appendx C. We run all methods to a stable convergence and measure the statistical error. The top row of Figure 1 shows the performance of these methods under different impact of statistical heterogeneity.
From Figure 1(a), in a controlled convex setting, we observe that FedProx consistently interpolates between GlobalTrain and LocalTrain. As $\lambda$ increases, FedProx transitions from behaving similarly to LocalTrain towards GlobalTrain. When statistical heterogeneity is low, both GlobalTrain and FedProx with larger $\lambda$ values achieve better statistical performance. However, as heterogeneity increases, LocalTrain and FedProx with smaller $\lambda$ values show better results. The figure also demonstrates that with an adaptive choice of $\lambda$ in theoretical analysis outlined in Theorem 1, FedProx consistently achieves a low error across varying levels of heterogeneity. See more results and discussion in Appendix C. Note that as statistical heterogeneity increases, LocalTrain experiences a slight rise in error due to the increased difficulty of the statistical problem. However, it remains the least affected by heterogeneity since it relies solely on local data. In Figure 1(b) and 1(c), for the MNIST dataset, which includes both convex logistic regression and non-convex CNN settings, we observe similar patterns. FedProx again strikes a balance between GlobalTrain and LocalTrain, adapting its performance as $\lambda$ changes. While GlobalTrain performs best in this scenario due to the relatively weak statistical heterogeneity, FedProx, with increasing $\lambda$, moves closer to GlobalTrain performance. In Figure 4 in Appendix C, we also evaluate the local model’s statistical error and observe that as statistical heterogeneity increases, FedProx with smaller $\lambda$ values can surpass GlobalTrain.
Figure 2: Effect of local updates on FedProx performance. Smaller $\lambda$ reduces communication with more local updates, while larger $\lambda$ benefits marginally from increasing local updates and requires more frequent communication.
6.3 Optimization Aspect: Communication-Computation Trade-off
In this section, we explore the trade-off between communication and computation in FedProx, controlled by varying the personalization degree.
Communication Cost. In the bottom row of Figure 1, we show the change in the total error of the algorithm (i.e., the difference of the model in $t$th round between the true underlying model) over communication rounds. The first observation is that as $\lambda$ increases, the convergence rate w.r.t. communication rounds slows across all settings. FedProx with a large $\lambda$ (closer to GlobalTrain) needs the most communication rounds to converge, while smaller $\lambda$ values (closer to LocalTrain) result in faster convergence. This result aligns with our theory (c.f. Corollary 1), showing that Fedprox with less personalization relies more heavily on communication. On the statistical side, Figure 1(d) shows that when statistical heterogeneity is low, larger $\lambda$ leads to the smallest total error. As $R$ increases, smaller $\lambda$ (more personalization) results in better performance.
Computation Cost. In Figure 2, we investigate how the number of local updates per communication round impacts the performance of FedProx at different personalization degree. Additional results in the non-convex setting can be found in Appendix C. Notice that for models with smaller $\lambda$ (higher personalization), increasing the number of local updates significantly improves convergence, making the algorithm less dependent on frequent communication. This aligns with our analysis: more personalized models rely less on communication and benefit more from local computation.

### v4

@@ 4.2 Effect of Personalization on Communication Efficiency
(ii) # gradient $\begin{split}&\text{(ii) \# gradient evaluations}=\tilde{\mathcal{O}}\Big{(}\kappa\cdot\log\frac{1}{\varepsilon}\Big{)}.\end{split}$
Corollary 2 clearly shows the influence of the personalization degree $\lambda$ on the communication cost. Specifically, as $\lambda$ increases from 00 to $\infty$, the communication complexity increases from $\mathcal{O}(\log(1/\epsilon))$ to $\mathcal{O}(\kappa\log(1/\epsilon))$. This fact corroborates our intuition that a higher degree of collaboration requires more communication resources. Moreover, the result given by Corollary 2 also indicates that the total computation cost is independent of $\lambda$. As such, we can see that model personalization can provably reduce the communication cost without any extra computation overhead. Notice that although personalization indeed enhances communication efficiency, doing so may deteriorate the statistical accuracy of the local models. In the next section, we will discuss in detail the trade-off between communication and statistical accuracy.
We also provide an extension of Alg. 1 to the stochastic setting using mini-batch stochastic gradients for the local updates. The proposed algorithm, termed FedCLUP, is given in Appendix B.5. Following similar argument to Theorem 2, Theorem 3 in Appendix B.5 establishes the convergence rate of FedCLUP, showing that it takes $\mathcal{O}\left(\kappa\frac{\lambda+\mu}{\lambda+L}\cdot\log\frac{1}{\varepsilon}\right)$ rounds of communication to obtain an output ${\bm{w}}_{t}^{(g)}$ within an error ball of size $\left(\mathcal{O}(\varepsilon)+\frac{\sigma^{2}}{\mu^{2}Bm}\right)$ around $\widetilde{\bm{w}}^{(g)}$. Here $\sigma^{2}$ is the variance of the stochastic gradient defined in Assumption 4 in Appendix B.5 and $B$ is the mini-batch size. Therefore, the influence of $\lambda$ on communication complexity is consistent with the noiseless setting presented in Corollary 2. Additionally, one may note that in the stochastic setting, the error due to the stochastic noise scales inversely with the number of participating clients, showing the advantage of client collaboration in reducing the variance of ${\bm{w}}_{t}^{(g)}$.
5 Communication-Accuracy Trade-off
Building on the results established in Section 4.1 and 4.2, we cast insights on the trade-off between communication efficiency and statistical accuracy, and discuss its implications on real-world practice.
Corollary 3.
Under the conditions of Theorem 1 and Corollary 2, ${\bm{w}}_{T,K}^{(i)}$ generated by Algorithm 1 satisfies
$\begin{split}&\mathbb{E}\left\|{\bm{w}}_{T,K}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\right\|^{2}\\ &\leq 2\left(1-\frac{1}{4\kappa}-\frac{L-\mu}{4\kappa(\lambda+\mu)}\right)^{T}\mathbb{E}\left\|{\bm{w}}_{0,0}^{(i)}-\widetilde{\bm{w}}^{(i)}\right\|^{2}\\ &+\mathcal{O}\left[\left(\frac{1}{N}+R^{2}+\frac{1}{q_{1}(\lambda)}\right)\wedge\left(\frac{1}{n}+q_{2}(\lambda)\right)\right]\end{split}$ (15)
for any $i\in[m]$, where $q_{1}(\lambda)$ and $q_{2}(\lambda)$, defined in (70) and (67), respectively, are monotonically increasing functions of $\lambda$ with $\lim_{\lambda\to\infty}q_{1}(\lambda)=\infty$ and $\lim_{\lambda\to 0}q_{2}(\lambda)=0$.
Methods Convexity Bounded Domain Computation Cost per Round Communication Cost Statistical Error
FedProx (Li et al., 2020) cvx (1) $\mathcal{O}(\Delta/(\rho\varepsilon))$
pFedMe (T Dinh et al., 2020) s-cvx (2) $\mathcal{O}(1/{(mR\varepsilon)})$
L2GD (Hanzely & Richtárik, 2020) s-cvx $\mathcal{O}(1+\frac{L}{\lambda})$ $\mathcal{O}({\frac{2\lambda\kappa}{\lambda+L}\log(\frac{1}{\varepsilon})})$
AL2SGD (Hanzely et al., 2020) s-cvx $\mathcal{O}\left(\frac{(m+\sqrt{m(L+\lambda)/\mu})}{\sqrt{\min\{L,\lambda\}/\mu}}\right)$ $\mathcal{O}(\sqrt{\min\{L,\lambda\}/\mu}\log\frac{1}{\varepsilon})$
Algorithm 3 (Chen et al., 2023c) s-cvx $\mathcal{O}(\lambda/\varepsilon)$ $\mathcal{O}((\lambda\vee 1)/\varepsilon)$
FedCLUP s-cvx $\mathcal{O}(\frac{\lambda+L}{\lambda+\mu})$ $\mathcal{O}(\kappa\frac{\lambda+\mu}{\lambda+L}\log{\frac{1}{\varepsilon}})$
@@ 5 Communication-Accuracy Trade-off
Table 1: Comparison of the existing works analyzing Problem (2) (cvx denotes convexity and s-cvx denotes strong convexity).
See Appendix C for the proof. When collaborative learning is beneficial, i.e., when $R^{2}\lesssim 1/n$, in this case, Corollary 1 establishes that the minimax-optimal statistical rate is $\mathcal{O}(1/N+R^{2})$. As $\lambda$ increases, $[q_{1}(\lambda)]^{-1}$ monotonically decreases to zero, leading to improved statistical accuracy, approaching the optimal accuracy. However, since the optimization error in (15) is monotonically increasing with $\lambda$, a higher degree of collaboration will also results in slower convergence over communication rounds. Consequently, Corollary 3 implies that when the data heterogeneity is relatively low, increasing personalization will improve communication efficiency but at the expense of lower statistical accuracy, and vice versa. Such an opposite effect of the personalization degree on the statistical accuracy and communication efficiency leads to an accuracy-communication trade-off.
As the optimization error vanishes when $T\to\infty$, the statistical error dominates the total error in Corollary 3. However, the associated communication cost becomes prohibitively large to achieve such an error. Therefore, balancing the optimization and statistical errors is crucial for achieving a target total error with efficient resource usage. Under the influence of the personalization degree, a practical guideline emerges: when collaborative learning is beneficial (i.e., $R^{2}\leq 1/\sqrt{n}$), to achieve a target magnitude of total error, we control the optimization errors at the same magnitude as the statistical error. Specifically, we increase $\lambda$ to reduce the statistical error to the same magnitude of the total error, and then gradually increase the communication rounds $T$ until the optimization error reaches the same magnitude as well. Since the communication efficiency will decrease will an increasing $\lambda$, the above strategy can achieve a given target total error at the highest possible communication efficiency.
Later, we will also show that, with the optimal choice of personalization degree, FedCLUP improves communication efficiency significantly compared with GlobalTrain and achieve a total error smaller than LocalTrain in Section 6.
Comparison with Existing Literature. As reported in Table 1, our work is the first to quantitatively characterize how changing the personalization degree leads to the trade-off between communication cost and statistical accuracy. In contrast, most prior studies either do not explicitly analyze statistical convergence or fail to provide a tight convergence guarantee for optimization and statistical error. For example, Chen et al. (2023c) attempted to establish both optimization and statistical rates simultaneously; however, when $\lambda\to\infty$, Problem (2) reduces to GlobalTrain, yet their results incorrectly imply that the communication and computation costs diverge to infinity.
6 Empirical Study
In this section, we provide empirical validation for our theoretical results, evaluating Problem (2) on both synthetic and real datasets with convex and non-convex loss functions. We first present the experimental setup, and then analyze the impact of personalization on statistical accuracy and communication efficiency, and conclude by analyzing the trade-off between the two. Details about the experimental setup are available in Section 6.1 and Appendix D, and the complete anonymized codebase is accessible at https://github.com/ZLHe0/fedclup.
@@ 6.1 Experimental Details
Synthetic Dataset. We generate two cases of synthetic dataset: (1)As our theoretical analysis is established under strong convexity, first, we consider an overdetermined linear regression task, where the choice of hyparamter is strictly followed Corollary 3. (2) We follow a similar procedure to prior works (Li et al., 2020) but with some modifications to align with the setup in this paper. Specifically, for each client we generate samples $(\bm{X}_{k},\bm{y}_{k})$, where the labels $\bm{y}_{k}$ are produced by a logistic regression model $\bm{y}_{k}=\text{argmax}(\sigma({\bm{w}}_{k}^{\top}\bm{X}_{k}))$, with $\sigma$ being the sigmoid function. The feature vectors $\bm{X}_{k}$ are drawn from a multivariate normal distribution $\mathcal{N}(\bm{v}_{k},\bm{\Sigma})$, where $v_{k}\sim\mathcal{N}(0,1)$ and the covariance matrix $\bm{\Sigma}$ follows a diagonal structure $\bm{\Sigma}_{j,j}=j^{-1.2}$. The heterogeneity is introduced by sampling the model weights ${\bm{w}}_{k}$ for each client from a normal distribution $\mathcal{N}(0,R)$, where $R\in[0,3]$ controls the statistical heterogeneity across clients’ data.
Dataset LocalTrain FedCLUP pFedMe GlobalTrain
High Medium Low High Medium Low
MNIST Logit 0.7828 0.8123 0.8213 0.8312 0.8001 0.8239 0.8291 0.8391
MNIST CNN2 0.8194 0.8753 0.8893 0.9032 0.8756 0.8771 0.8749 0.9098
MNIST CNN5 0.7633 0.9091 0.9102 0.9421 0.9340 0.9385 0.9440 0.9433
EMNIST CNN5 0.6111 0.6278 0.6415 0.6672 0.6345 0.6532 0.6717 0.6611
CIFAR10 CNN3 0.6238 0.6102 0.6712 0.7302 0.6744 0.7443 0.7732 0.8041
Table 2: Test accuracy of algorithms across datasets under varying personalization degrees. “Logit” refers to logistic regression, while “CNN-k” denotes a convolutional neural network with k convolutional layers. “Low,” “Medium,” and “High” indicate low, moderate, and high personalization levels, respectively. As the personalization degree decreases, the statistical accuracy of FedCLUP and pFedMe transitions from resembling LocalTrain to GlobalTrain.
Real Dataset. We use the MNIST, EMNIST, CIFAR10, Sent140 an CelebA datasets for real data analysis. For MNIST, EMNIST and CIFAR10, they aren’t naturally partitioned datasets. Therefore, following Li et al. (2020), to impose statistical heterogeneity, we distribute the data across clients in a way that each client only has access to a fixed number of classes. The fewer classes each client has access to, the higher the statistical heterogeneity. Details on data preprocessing and heterogeneity settings for each dataset are provided in Appendix D. For Sent140 and CelebA datasets, as they are naturally partitioned, we don’t need to create the data heterogeneity and directly apply our algorithm to these datasets with advanced models, like CNN and LSTM model (Sak et al., 2014). We strictly follow the previous work (Li et al., 2021) and (Duan et al., 2021) to pre-process the Sent140 and CelebA dataset, respectively.
@@
Implementation and Evaluation. For each setup, we evaluate FedCLUP with three different degree of personalization: low, medium, and high, with specific personalization degree varying based on the dataset (details provided in Appendix D). For the synthetic dataset, in alignment with theorem 2, we implement FedCLUP using a global step size $\gamma=(\lambda+L)/(\lambda L)$ and a local step size $\eta=(L+\lambda)^{-1}$. For the real dataset, since $L$ is unknown, we implement the same algorithm with a global step size $\gamma$ set to $1/\lambda$, while the local step size is determined via grid search. In terms of evaluation, for the synthetic dataset, we evaluate by tracking the ground truth models and measuring error as the distance from these ground true models. For the real datasets, we report training loss and testing accuracy. Additional details, including hyperparameter settings and evaluation metric definition, can be found in Appendix D.
Figure 1: Total error and training loss of FedCLUP with varying personalization degrees. Total error (y-axis in (a)) quantifies the distance between the model at each communication round and the ground truth model, i.e. $\|{\bm{w}}_{T,K}^{(i)}-{{\bm{w}}_{\star}^{(i)}}\|^{2}$. (a) compares total error under low ($R=0.5$) and high ($R=2.0$) heterogeneity. (b) presents training loss across different datasets and models under low heterogeneity. Synthetic-Logit and MNIST-Logit represent logistic regression on synthetic and MNIST data, respectively, while MNIST-CNN, CIFAR10-CNN, and EMNIST-CNN represent CNN models trained on MNIST, CIFAR-10, and EMNIST datasets. In a low-heterogeneity setting, higher personalization (high per) accelerates convergence, while lower personalization (low per) improves final error. In a high-heterogeneity setting, higher personalization achieves both smaller error and faster convergence.
6.2 Results
Effect of Personalization on Statistical Accuracy. We compare FedCLUP, under different personalization degrees, with GlobalTrain, LocalTrain and pFedMe (T Dinh et al., 2020), an alternative algorithm for solving Problem (2). Implementation details are provided in Appendix D. All methods are run until stable convergence, and their test accuracy is reported in Table 2 in the low heterogeneity setting. Table 2 shows that as the personalization degree decreases and collaborative learning increases, the solution of Problem (2) becomes closer to GlobalTrain, leading to improved statistical accuracy due to increased information sharing across clients. A similar trend is observed for both FedCLUP and pFedMe across different datasets and models, showing that under low heterogeneity, increased collaboration improves statistical accuracy.
Figure 2: Effect of personalization on communication efficiency in FedCLUP for Synthetic-Logit (left) and MNIST-Logit (right). Each subfigure compares the impact of varying the number of local updates per communication round under different personalization levels.
Effect of Personalization on Communication Efficiency. Figure 2 investigates how personalization impacts communication efficiency in FedCLUP by analyzing the benefit of increasing local updates under different personalization levels. In the low-personalization setting (left column), more local updates significantly accelerate convergence, reducing the reliance on frequent communication. However, in the high-personalization setting (right column), increasing local updates has a limited effect on convergence, indicating that frequent communication is essential for effective learning. This demonstrates that higher personalization requires more communication rounds to achieve comparable performance, aligning with our theoretical findings, showing that increased personalization improves communication efficiency.
The Trade-off under Personalization. The above two observations imply that there shall be a accuracy-communication trade-off in choosing the personalization degree to achieve the smallest total error under given communication budget (See Figure 1). To further demonstrate this point, in the left part of Figure 3, we compare the total error over iterations among LocalTrain, GlobalTrain, and FedCLUP under varying levels of personalization. Once the algorithms converge, for any given total error, there exists a specific degree of personalization in FedCLUP that incurs the least communication cost among all considered methods. This observation supports the conclusion in Corollary 3 that a unique personalization degree exists in FedCLUP to achieve communication efficiency for a fixed total error. Furthermore, we observe that no fixed personalization degree consistently outperforms others across all error levels. This implies that adaptively adjusting the personalization degree is essential for FedCLUP to maintain communication efficiency across different total error regimes. To further validate this finding, we conduct the same experiment on real-world datasets. Results on the CelebA and Sent140 datasets are presented in Figures 9 and 10, respectively. Additionally, we observe a similar phenomenon in pFedMe, another method solving Problem 2, when we vary its personalization degree. Witnessing the insightful results, it motivates us to adaptively change the personalization degree in FedCLUP for achieving communication efficiency over different total errors.
Potential Solution in Practice. After understanding the trade-off under personalization, we propose a dynamic tuning strategy for the optimal $\lambda$ with communication efficiency(See Figure 3 Right ). Beginning with a small $\lambda$ for efficiency, as soon as the validation performance plateaus or the statistical error stops improving significantly, we gradually increase $\lambda$. This allows the model to benefit from enhanced generalization through increased collaboration at the expense of slightly higher communication costs. By progressively adjusting in this way, one can finally identify the optimal personalization degree that balances the trade-off and minimizes the total communication cost required to meet a desired performance threshold.
Figure 3: Total error versus communication rounds. Left: Comparison among LocalTrain, GlobalTrain, and FedCLUP with varying levels of personalization. For a given target total error, an optimal $\lambda$ exists that has the minimum communication rounds. Right: A dynamic $\lambda$ strategy: we begin with local training (highest personalization degree), and once the decrease of the validation error slows down, we gradually decrease the personalization degree. Comparing the right figure with the left one shows that the dynamic strategy approximates the optimal personalization level needed to meet a desired error threshold with minimal communication rounds, demonstrating the practical effectiveness of tuning personalization dynamically according to the accuracy-communication trade-off.